### 04 Mar 2026 v1

# Resolving Spurious Multifractality in Discrete Systems: A Finite-Size Scaling Protocol for MFDFA in the 2D Ising Model

Abstract
Multifractal Detrended Fluctuation Analysis (MFDFA) has emerged as a standard tool for characterizing scale invariance in complex systems, yet its application to discrete spin models is frequently marred by reports of “spurious multifractality” that contradict established theory. In this work, we resolve this controversy by establishing a rigorous protocol for the analysis of discrete lattice snapshots. Using the 2D Ising model as a benchmark, we demonstrate that the previously reported broad singularity spectra [1] are finite-size artifacts dominated by lattice discreteness effects in the negative moment regime ($q<0$). By restricting the analysis to positive moments and performing a systematic Finite-Size Scaling (FSS) analysis, we show that the spectral width collapses to zero ($\Delta\alpha\to 0$) in the thermodynamic limit. The method accurately recovers the monofractal exponent of the Ising universality class ($\alpha\approx H\approx 0.875$), consistent with Conformal Field Theory. To validate the discriminatory power of this protocol, we contrast these findings with the Random Bond Ising Model (RBIM), showing that quenched disorder induces a genuine, broad multifractal spectrum ($\Delta\alpha\approx 0.23$) that survives scaling. Furthermore, we propose a theoretical interpretation where the MFDFA polynomial detrending functions as a phenomenological Renormalization Group filter, suppressing analytic background fields (irrelevant operators) to isolate the singular critical behavior. These results define a robust methodology for distinguishing between clean and disorder-dominated criticality in finite systems.
I Introduction
The two-dimensional (2D) Ising model stands as a cornerstone of statistical mechanics, serving as the prototypical system for understanding continuous phase transitions and universality. Since Onsager’s exact solution in the thermodynamic limit [2], the model has provided a rigorous testing ground for the development of Renormalization Group (RG) theory and Conformal Field Theory (CFT)[3]. It is well-established that near the critical temperature $T_{c}\approx 2.269$ ($J/k_{B}$), the system exhibits scale invariance governed by the 2D Ising universality class. Specifically, the spin-spin correlation function decays algebraically as:
$G(r)=\langle\sigma_{i}\sigma_{i+r}\rangle\sim\frac{1}{r^{d-2+\eta}},$ (1)
where $d=2$ is the spatial dimension and $\eta=1/4$ is the anomalous dimension [3]. A fundamental prediction of this theoretical framework is that the critical fixed point is strictly monofractal. That is, the scaling is uniform across the lattice, governed by a single relevant thermal operator, and the singularity spectrum $f(\alpha)$ collapses to a single point.
Despite this theoretical clarity, the characterization of critical fluctuations via modern signal processing techniques has proven surprisingly contentious. In recent decades, Multifractal Detrended Fluctuation Analysis (MFDFA) [4, 5] has emerged as a powerful tool for quantifying the scaling properties of non-stationary signals. Originally developed for complex time series in fields ranging from turbulence and geophysics to quantitative finance, MFDFA generalizes standard fluctuation analysis by extracting a full spectrum of generalized Hurst exponents $h(q)$. While the method has demonstrated remarkable success in continuous systems, its application to discrete spatial structures—such as spin lattice snapshots—has generated conflicting results.
The importance of resolving this controversy extends beyond the 2D Ising model itself. MFDFA has become a standard tool across disciplines—from econophysics and geophysics to biophysics and climate science —where practitioners often apply the method to discrete or discretized data without awareness of the potential pitfalls identified here. Our systematic resolution provides a validated protocol that will enable researchers in these fields to distinguish genuine multifractal signatures from finite-size artifacts, thereby enhancing the reliability of multifractal characterization across the physical sciences.
Moreover, this work demonstrates a broader methodological principle: signal processing techniques developed for continuous systems (turbulence, financial time series) require careful adaptation when applied to discrete lattice models. The different physical meaning of “small fluctuations” in continuous versus discrete systems necessitates moment-range restrictions that would be unnecessary in the original application domains.
A survey of the recent literature reveals a persistent anomaly: numerous numerical studies report broad multifractal spectra in the pure 2D Ising model. These findings imply the existence of a hierarchy of scaling exponents, a result that contradicts the unique scaling dimension mandated by the RG fixed point. This discrepancy—often termed ”spurious multifractality”—raises a fundamental question: Does the reported multifractality reflect hidden complexity in the Ising phase transition, or does it signal a breakdown of the MFDFA methodology when applied to discrete variables in finite systems?
In this work, we resolve this controversy by establishing a rigorous protocol for the multifractal analysis of discrete lattice systems. We demonstrate that the apparent multifractality is not an intrinsic property of the Ising universality class, but a methodological artifact arising from the interplay between finite system sizes and lattice discreteness. Specifically, we identify that the divergence of the singularity spectrum is driven largely by negative moments ($q<0$). In continuous fields (e.g., fluid turbulence), negative moments probe rare, small-amplitude fluctuations. However, in discrete spin models, these moments probe ”frozen” domains where the local variance vanishes due to the lattice cutoff.
By restricting the analysis to the physically relevant regime of positive moments ($q>0$) and performing a systematic Finite-Size Scaling (FSS) analysis, we show that the spectral width $\Delta\alpha$ scales as a power law of the system size, $L$, collapsing to zero in the thermodynamic limit ($L\to\infty$). The method accurately recovers the monofractal Hurst exponent $H\approx 0.875$, in precise agreement with the CFT prediction $H=1-\eta/2$, where $\eta=1/4$ is the anomalous dimension.
To validate that this collapse to monofractality is a correct physical result rather than a loss of methodological sensitivity, we contrast our findings with the Random Bond Ising Model (RBIM). The RBIM, characterized by quenched bond disorder, is known to exhibit ”Griffiths phases” [6] —rare regions that order locally at temperatures different from the bulk $T_{c}$. We demonstrate that our MFDFA protocol successfully detects a genuine, broad multifractal spectrum ($\Delta\alpha\approx 0.23$) in the disordered case that survives finite-size scaling. This establishes the method as a robust discriminator between ”clean” criticality and ”dirty” multifractality.
Furthermore, we propose a theoretical reinterpretation of the MFDFA procedure. We argue that the polynomial detrending step acts as a phenomenological projection operator in the RG sense. By filtering out analytic background trends—which correspond to irrelevant operators in the Hamiltonian—MFDFA effectively isolates the singular scaling part of the free energy density.
The remainder of this paper is organized as follows. Section II outlines the theoretical connection between fluctuation analysis and critical exponents. Section III details the numerical protocol, emphasizing the treatment of discrete limits. Section IV presents the Finite-Size Scaling results for the pure model and the comparative analysis of the RBIM. Finally, Section VI discusses the implications of these findings for the analysis of experimental data where the underlying universality class may be unknown.
II Theoretical Framework: MFDFA as a Real-Space Renormalization Transform
While MFDFA is primarily a signal processing tool, its application to critical phenomena admits a rigorous physical interpretation. We propose that the MFDFA procedure acts as an empirical implementation of a Real-Space Renormalization Group (RG) transformation, where the detrending step plays the role of filtering out irrelevant operators.
II.1 Fluctuation Function and Correlation Scaling
In the context of the 2D Ising model, the integrated profile $Y(i)$ analyzed by MFDFA corresponds to the coarse-grained magnetization field. For a 1D spatial cut of the lattice, the profile at position $x$ is defined as the running sum of the spin variables $\sigma(x^{\prime})$:
$Y(x)=\sum_{i=1}^{x}\sigma_{i}\approx\int_{0}^{x}\sigma(x^{\prime})dx^{\prime}$ (2)
The central quantity in MFDFA is the local variance $F^{2}(s)$ within a window of size $s$. Physically, this variance measures the fluctuations of the coarse-grained magnetization operator $\hat{M}_{s}$ at scale $s$. Neglecting the detrending for a moment, the second-order fluctuation function $F_{2}^{2}(s)$ is directly related to the two-point correlation function $G(r)=\langle\sigma(x)\sigma(x+r)\rangle$:
$F_{2}^{2}(s)\sim\langle M_{s}^{2}\rangle=\int_{0}^{s}dr\int_{0}^{s}dr^{\prime}\,G(|r-r^{\prime}|)$ (3)
Near the critical temperature $T_{c}$, substituting the power-law decay Eq. (1) and integrating twice yields the scaling of the fluctuation function:
$F_{2}^{2}(s)\sim s^{2-(d-2+\eta)}=s^{4-d-\eta}$ (4)
Since $F_{2}(s)\sim s^{2H}$, we obtain the fundamental relation connecting the Hurst exponent with the critical exponents of the universality class. For a 2D bulk system ($d=2$), this simplifies to:
$H=1-\frac{\eta}{2}$ (5)
Substituting the exact value $\eta=0.25$, we obtain a precise theoretical prediction for the critical point:
$H(T_{c})=1-\frac{0.25}{2}=0.875$ (6)
II.2 Detrending as an Irrelevant Operator Filter
A key feature of MFDFA is the subtraction of a polynomial trend $P_{\nu}(x)$ of order $\nu$ from the integrated profile. While often treated as a mere baseline correction, this step admits a physical interpretation in the context of critical phenomena, as illustrated in Figure 1.
[…]
Figure 1: Schematic of the MFDFA detrending interpreted as an RG filter. The raw integrated profile $Y(x)$ (top, blue) is a superposition of singular critical fluctuations and a smooth analytic background. The polynomial trend $P_{\nu}(x)$ (red dashed) captures the macroscopic background field driven by irrelevant operators. The MFDFA detrending operation acts as a projection $\hat{\mathcal{P}}_{\nu}$, subtracting this background to isolate the singular scaling part (bottom, green). This filtering grants access to the intrinsic critical exponents even in finite systems where correction-to-scaling terms are significant.
The local magnetization density $m(x)$ near a critical point can be decomposed into a singular scaling part $m_{sing}(x)$ (governed by the relevant thermal operator) and a regular analytic part $m_{reg}(x)$ (arising from short-range fluctuations and irrelevant operators):
$m(x)=m_{sing}(x)+m_{reg}(x).$ (7)
The regular part $m_{reg}(x)$ is smooth and analytic, effectively creating a ”background” trend that obscures the fractal scaling of $m_{sing}(x)$ in finite systems. The MFDFA detrending operation, $Y(x)\to Y(x)-P_{\nu}(x)$, functions as a local projection operator $\hat{\mathcal{P}}_{\nu}$. Since $P_{\nu}(x)$ captures smooth variations up to order $\nu$, it effectively filters out the contributions of $m_{reg}(x)$. Consequently, the fluctuation function $F_{q}(s)$ primarily probes the singular component:
$F_{q}(s)\sim\langle|\hat{\mathcal{P}}_{\nu}[Y_{sing}]|^{q}\rangle^{1/q}\sim s^{\tau(q)/q}.$ (8)
By suppressing these analytic corrections to scaling, MFDFA allows for the extraction of asymptotic critical exponents even from system sizes where irrelevant operators would typically distort the signal.
III Numerical Protocol
Our investigation combines extensive Monte Carlo simulations with a rigorous application of the MFDFA formalism adapted for discrete spatial structures. In this section, we detail the lattice models employed and the specific signal processing pipeline used to extract scaling exponents.
III.1 Model Hamiltonians and Monte Carlo Scheme
We consider the 2D Ising model on a square lattice of linear dimension $L$, with periodic boundary conditions. The system is described by the Hamiltonian:
$\mathcal{H}=-\sum_{\langle i,j\rangle}J_{ij}\sigma_{i}\sigma_{j},$ (9)
where $\sigma_{i}\in\{-1,+1\}$ represents the spin variable at site $i$, and the sum runs over nearest-neighbor pairs. For the pure Ising model, the ferromagnetic coupling is uniform, $J_{ij}=J>0$, leading to a critical point at $T_{c}\approx 2.269J/k_{B}$. To investigate the effects of quenched disorder, we also simulate the Random Bond Ising Model (RBIM), where the interactions $J_{ij}$ are drawn from a static probability distribution. This introduces local impurities that pin domain walls and generate ”Griffiths phases”—rare regions that order locally at temperatures deviating from the bulk critical point.
The equilibrium configurations were generated using the Metropolis single-spin flip algorithm. To accommodate the critical slowing down near $T_{c}$, the number of Monte Carlo steps (MCS) per site was dynamically adjusted with system size, typically scaling as $L^{z}$ with $z\approx 2.17$. We simulated lattice sizes ranging from $L=32$ to $L=256$. For each temperature and system size, we collected an ensemble of $10^{4}$ independent equilibrium snapshots to ensure statistical convergence of the high-order moments.
III.2 Multifractal Analysis of Spatial Profiles
To characterize the scale-invariant properties of the spin configurations, we employ the Multifractal Detrended Fluctuation Analysis (MFDFA) [4]. This method generalizes standard fluctuation analysis to a full spectrum of moments, making it particularly suitable for detecting intermittent scaling in spatially heterogeneous media. We employ MFDFA rather than wavelet-based alternatives (e.g., WTMM) because comparative studies have shown MFDFA to be more numerically stable for identifying strict monofractality in short time series [7], making it ideal for finite-size scaling analysis.
For completeness, we provide below a concise description of the standard MFDFA algorithm as adapted for discrete spatial profiles.
1. Profile Construction:
We first map the 2D lattice snapshots into 1D spatial profiles by defining a geometric cut along the $x$-direction. The input series corresponds to the sequence of spins $\{\sigma_{i}\}_{i=1}^{L}$ along this cut. We construct the integrated profile $Y(k)$ by accumulating the fluctuations around the global magnetization:
$Y(k)=\sum_{i=1}^{k}(\sigma_{i}-\langle\sigma\rangle),\quad k=1,\dots,L$ (10)
where $\langle\sigma\rangle$ is the ensemble average magnetization.
2. Segmentation and Detrending:
The profile $Y(k)$ is divided into $N_{s}=\lfloor L/s\rfloor$ non-overlapping segments of scale $s$. To reduce data loss at the boundaries, we repeat the segmentation starting from the opposite end, yielding $2N_{s}$ segments. In each segment $v$, we calculate the local trend by fitting a polynomial $P_{\nu}^{(v)}$ of order $\nu$ (typically $\nu=1$ or $2$) to the data. The detrended variance is then computed as:
$F^{2}(s,v)=\frac{1}{s}\sum_{j=1}^{s}\{Y[(v-1)s+j]-P_{\nu}^{(v)}(j)\}^{2}.$ (11)
As discussed in Sec. II, this detrending step acts as a high-pass filter, removing analytic background fields associated with irrelevant operators.
3. Fluctuation Function:
We obtain the $q$-th order fluctuation function by averaging over all segments:
$F_{q}(s)=\left\{\frac{1}{2N_{s}}\sum_{v=1}^{2N_{s}}[F^{2}(s,v)]^{q/2}\right\}^{1/q},\quad\text{for }q\neq 0.$ (12)
For the special case $q\to 0$, the averaging procedure is given by the exponential of the logarithmic mean:
$F_{0}(s)=\exp\left\{\frac{1}{4N_{s}}\sum_{v=1}^{2N_{s}}\ln[F^{2}(s,v)]\right\}.$ (13)
4. Scaling and Singularity Spectrum:
The scaling behavior is determined by the power-law relation $F_{q}(s)\sim s^{h(q)}$, where $h(q)$ is the generalized Hurst exponent. Positive moments ($q>0$) emphasize large fluctuations (critical clusters), while negative moments ($q<0$) amplify small fluctuations. From $h(q)$, we derive the mass exponent $\tau(q)$:
$\tau(q)=qh(q)-1.$ (14)
Finally, the multifractal singularity spectrum $f(\alpha)$ is obtained via the Legendre transform:
$\alpha(q)$ $=\frac{d\tau(q)}{dq},$ (15)
$f(\alpha)$ $=q\alpha(q)-\tau(q).$ (16)
Here, $\alpha$ represents the local Hölder exponent (singularity strength), and $f(\alpha)$ denotes the fractal dimension of the subset of the series characterized by $\alpha$. A monofractal system (like the pure Ising model) is characterized by a linear $\tau(q)$ and a singularity spectrum that collapses to a single point, $f(\alpha)\to 1$ at $\alpha=H$.
The MFDFA implementation employed here follows the standard algorithm established by Kantelhardt et al. [4], with optimizations for computational efficiency when analyzing large ensembles of lattice configurations. Our research group has successfully applied MFDFA and related multifractal methods to diverse systems including climate indices [8], financial time series [9], and biological sequences [10], establishing the methodological foundation for the present investigation of discrete lattice models.
III.3 The Discrete Limit and Physical Cutoffs ($q>0$)
Special care must be taken regarding the discrete nature of Ising spins ($\sigma\in\{-1,1\}$). Standard multifractal formalism includes negative moments ($q<0$), which weigh small fluctuations heavily. However, in discrete lattice models, the limit of ”small fluctuations” is fundamentally different from that in continuous fields (e.g., turbulence). In the Ising model, a ”small fluctuation” region often corresponds to a ”frozen” domain where the local variance is exactly zero (or strictly limited by the lattice spacing $a$). As the box size $s$ decreases, the variance $F^{2}(s)$ in these regions hits a hard physical floor determined by the discreteness of the spin variables, rather than a continuous dissipation scale. Consequently, the scaling of negative moments is dominated by the lattice cutoff rather than the correlation length $\xi$. To avoid mixing these lattice artifacts with genuine critical scaling, we restrict our analysis to the range of positive moments $q\in[0.5,5.0]$. In this regime, the partition function is dominated by large, active critical clusters, ensuring that the measured exponents reflect the universal continuum physics rather than the discreteness of the underlying grid.
IV Results
IV.1 The Critical Fixed Point: Restoration of Monofractality
We first examine the scaling behavior of the pure 2D Ising model. A prerequisite for any valid multifractal analysis is the accurate recovery of the known Hurst exponent at the critical point. Figure 2 displays the temperature dependence of the second-order Hurst exponent $H(q=2)$ for a large lattice ($L=256$). As the system approaches the critical temperature $T_{c}\approx 2.269$, the exponent exhibits a sharp crossover, intersecting the theoretical prediction $H=0.875$ precisely at the transition point. This agreement confirms that the spatial MFDFA protocol correctly captures the anomalous dimension of the magnetization field, satisfying the scaling relation $H=1-\eta/2$.
Figure 2: Evolution of the Hurst exponent $H$ ($q=2$) vs Temperature for $L=256$. The intersection with the theoretical value $H=0.875$ (blue dotted line) occurs precisely at the critical temperature $T_{c}$ (red dashed line).
The central result of this study concerns the width of the singularity spectrum $f(\alpha)$. Previous studies reporting ”spurious multifractality” typically observed a broad spectral width $\Delta\alpha=\alpha_{max}-\alpha_{min}>0$. In Figure 3, we present the evolution of the singularity spectrum for system sizes ranging from $L=32$ to $L=256$, restricted to the positive moment regime ($q>0$). For small system sizes ($L=32$), the spectrum indeed appears broad, mimicking a multifractal structure. However, this width is not intrinsic. As $L$ increases, we observe a dramatic narrowing of the spectrum. The distinct branches of the $f(\alpha)$ curve progressively collapse onto a single vertical line centered at $\alpha\approx 0.875$.
Figure 3: Evolution of the singularity spectra $f(\alpha)$ for $q>0$. Darker curves correspond to smaller system sizes ($L=32$), while lighter curves represent larger systems ($L=256$). A clear narrowing effect is observed as $L$ increases.
To quantify this convergence, we performed a Finite-Size Scaling (FSS) analysis of the spectral width. Figure 4 demonstrates that the average singularity exponent $\langle\alpha\rangle$ converges asymptotically to the theoretical Ising value:
$\lim_{L\to\infty}\langle\alpha\rangle_{L}=0.875\pm 0.002.$ (17)
Crucially, the spectral width $\Delta\alpha$ scales as a power law of the system size, $\Delta\alpha\sim L^{-\gamma}$, vanishing in the thermodynamic limit. This collapse provides strong evidence that the multifractality observed in the pure Ising model is most likely a finite-size artifact largely influenced by the mixing of irrelevant operators and discreteness effects at small scales.
Figure 4: Finite-Size Scaling (FSS) of the singularity exponent $\langle\alpha\rangle$. The convergence towards the theoretical value $\alpha=0.875$ (green line) is evident as the system size increases ($1/L\to 0$).
IV.2 Probing Disorder: The RBIM as a Control
To demonstrate that the observed collapse to monofractality is not due to a lack of sensitivity in our method, we contrast these findings with the Random Bond Ising Model (RBIM). The RBIM represents a distinct universality class where quenched disorder is relevant (by the Harris criterion [11]), leading to the formation of rare Griffiths phases that support a true spectrum of local correlation lengths.
Figure 5 compares the singularity spectrum of the pure model against the disordered model at the same system size. The contrast is stark:
  • Pure Model: The spectrum collapses to a vertical line ($\Delta\alpha\to 0$), indicative of a single scaling dimension.
  • Disordered Model (RBIM): The spectrum remains broad and robust ($\Delta\alpha\approx 0.23$), persisting even after finite-size scaling.
This persistence confirms that our MFDFA protocol is sensitive enough to detect genuine multifractality when it exists. The ”dirty” criticality of the RBIM generates a legitimate hierarchy of exponents due to the spatial heterogeneity of the bond disorder, whereas the ”clean” criticality of the pure model is strictly monofractal. Consequently, the MFDFA spectral width $\Delta\alpha$ act as an effective diagnostic indicator for distinguishing disorder-dominated phases from pure critical fixed points.
Figure 5: Sensitivity of MFDFA to disorder. While the pure Ising model converges to a vertical line (green dashed line, $\Delta\alpha\to 0$), the Random Bond Ising Model (RBIM, red circles) [12] displays a broad multifractal spectrum ($\Delta\alpha\approx 0.23$), revealing the complex fluctuation structure induced by quenched impurities.
V Methodological Robustness and Comparison with Alternative Approaches
Given the empirical nature of the Multifractal Detrended Fluctuation Analysis (MFDFA), it is essential to assess the robustness of the reported results with respect to methodological choices and to compare its performance against more conventional observables. In this section, we verify the stability of our results and discuss their relation to standard correlation functions.
V.1 Dependence on the Order of Detrending
[…]
To test the sensitivity of our results, we repeated the analysis at the critical temperature ($T\approx 2.27$) using DFA2 (quadratic detrending). We find that the estimated Hurst exponent improves in accuracy with higher-order detrending:
$H_{\mathrm{DFA1}}=0.826\pm 0.17,\qquad H_{\mathrm{DFA2}}=0.876\pm 0.13.$ (18)
Notably, the value obtained via quadratic detrending ($H\approx 0.876$) is in remarkably close agreement with the theoretical prediction ($H=0.875$). This suggests that the analytic background $m_{reg}(x)$ contains significant non-linear components that are better filtered by $P_{2}$ than $P_{1}$, validating our interpretation of detrending as a correction-to-scaling filter.
Although DFA2 offers superior numerical precision by more effectively filtering the non-linear analytic background, the central finding of this work, the collapse of the spectral width $\Delta\alpha$, is clearly evident even under linear detrending (DFA1), underscoring the robustness of the proposed protocol.
V.2 Stability with Respect to the Scaling Range
Another critical aspect is the selection of the scaling window $s\in[s_{\min},s_{\max}]$ used to extract the exponent. To evaluate stability, we systematically varied the fitting bounds for the $L=256$ system, testing ranges from short-scale focused ($s\in[8,40]$) to broad-scale ($s\in[12,80]$).
We observe that the extracted Hurst exponent is highly stable, with a relative variation of approximately $1.9\%$ across all tested ranges. This stability confirms the existence of a robust, scale-invariant regime at criticality, distinct from microscopic lattice artifacts or finite-size boundary effects.
V.3 Shuffling Tests
To further validate that the measured correlations are intrinsic to the critical spin configurations and not a trivial consequence of the marginal spin distribution, we performed shuffling tests. Specifically, we randomly shuffled the spin values along each one-dimensional cut, thereby destroying spatial correlations while preserving the single-site distribution $\sigma\in\{-1,1\}$.
For sufficiently large system sizes ($L\geq 128$) where the scaling range allows for robust fitting, the shuffled profiles consistently yield a Hurst exponent consistent with the random walk limit ($H=0.5$). For our largest system ($L=256$), we obtained:
$H_{\mathrm{shuffled}}=0.498\pm 0.07.$ (19)
The stark contrast between this null result and the critical value observed in the original data ($H\approx 0.88$) confirms that the MFDFA signal originates from genuine long-range spatial correlations and not from finite-amplitude effects or probability density artifacts.
V.4 Comparison with Two-Point Correlation Functions
The standard diagnostic of criticality in the Ising model is the two-point correlation function $G(r)=\langle\sigma(0)\sigma(r)\rangle$, which at criticality decays as $G(r)\sim r^{-\eta}$. While conceptually transparent, direct estimation of $G(r)$ in finite systems is often hampered by statistical noise at large distances, requiring extensive ensemble averaging to obtain clean power laws.
In contrast, MFDFA effectively integrates the correlation function twice (Eq. 4), which acts as a low-pass filter for the noise. The precise recovery of the theoretical exponent $H\approx 0.875$ (implying $\eta=0.25$) confirms that MFDFA captures the exact same scaling information as $G(r)$ but offers a more stable estimator in finite systems. The integration process smoothes the local fluctuations inherent to $G(r)$, allowing for a robust extraction of the anomalous dimension without the need for excessive averaging.
V.5 Relation to Cluster-Based Observables
An alternative real-space characterization of criticality is provided by cluster statistics, such as the distribution of Fortuin–Kasteleyn or geometric cluster sizes. These observables probe the fractal geometry of critical spin domains and are directly related to percolation theory.
While cluster analysis yields detailed geometric information, it typically requires nontrivial identification algorithms (e.g., Hoshen-Kopelman) and introduces additional definitions regarding connectivity. By contrast, MFDFA operates directly on raw spin configurations without any preprocessing. The recovery of the correct anomalous dimension $\eta$ demonstrates that MFDFA captures the same underlying scale invariance encoded in cluster observables, but in a computationally simpler and more flexible framework that requires no prior definition of ”clusters.”
V.6 Summary
The analyses presented in this section demonstrate that the critical scaling detected by our protocol is robust. The improvement in accuracy with DFA2, the stability across scaling ranges ($2\%$), and the null result for surrogates collectively validate MFDFA as a reliable probe for the 2D Ising phase transition, offering advantages in stability over direct correlation function measurements.
VI Discussion
Our investigation yields three interconnected insights that clarify the role of multifractal analysis in critical phenomena, resolving a long-standing ambiguity regarding the 2D Ising model.
VI.1 Resolution of the ”Spurious Multifractality” Controversy
The debate regarding the multifractal nature of the 2D Ising model has historically been fueled by analyses that indiscriminately include negative moments ($q<0$) or rely on small system sizes. Our Finite-Size Scaling results (Fig. 3 and Fig. 4) decisively show that the spectral width $\Delta\alpha$ is not an intrinsic property of the pure Ising fixed point, but a finite-size artifact.
We have identified the physical origin of this artifact: in discrete spin systems, negative moments probe the ”frozen” limit where local variance vanishes due to the lattice cutoff. Unlike continuous turbulent fields where small fluctuations scale universally, the ”small fluctuation” regime in a lattice model is dominated by discreteness effects. By restricting the analysis to positive moments ($q>0$) where critical clusters dominate, the system flows towards a strict monofractal limit defined by a single scaling dimension $\alpha=H\approx 0.875$. This aligns perfectly with the CFT prediction for the central charge $c=1/2$, confirming that only one relevant thermal operator governs the correlation decay.
VI.2 MFDFA as a Configural Diagnostic for Disorder
The contrast provided by the RBIM results (Fig. 5) is paramount for the validation of the method. If MFDFA were simply insensitive to subtle correlation structures, it would have failed to detect the complexity of the disordered model. The fact that the spectrum remains broad ($\Delta\alpha\approx 0.23$) for the RBIM in the same range of $L$ where the pure model collapses proves that MFDFA acts as a high-precision discriminator.
Physically, the RBIM is characterized by ”Griffiths phases”—rare regions devoid of impurities that order locally at temperatures different from the bulk $T_{c}$. These rare regions introduce a spatial distribution of local correlation lengths $\xi(\mathbf{r})$, which manifests mathematically as a true spectrum of singularities $f(\alpha)$. Therefore, we establish the spectral width $\Delta\alpha$ as an effective order parameter for distinguishing between ”clean” criticality (monofractal) and ”dirty” criticality (multifractal) in experimental data where the underlying Hamiltonian might be unknown.
VI.3 Renormalization Group Interpretation
Finally, the successful recovery of the anomalous dimension $\eta$ supports our theoretical proposal in Sec. II. The polynomial detrending $P_{\nu}(x)$ inherent to MFDFA should not be viewed merely as a baseline correction, but as a phenomenological projection operator in the RG space.
As illustrated in Fig. 1, the total free energy density contains both a singular critical part and a regular analytic background. The background arises from short-range fluctuations and irrelevant operators that have not yet decayed at the scale of observation. By removing these low-frequency analytic trends, MFDFA effectively ”renormalizes” the field configuration, projecting the signal onto the singular scaling manifold. This analogy suggests that varying the detrending order (DFA1, DFA2) could offer a systematic way to suppress higher-order correction-to-scaling terms, allowing for the extraction of asymptotic exponents from finite-size data with significantly reduced systematic error.
VI.4 Generality Beyond the 2D Ising Model
While we have focused on the 2D Ising model as the paradigmatic test case, the protocol developed here has broader applicability across lattice models and universality classes. The central insight—that negative moments in discrete systems probe lattice-scale discreteness rather than genuine critical fluctuations—is a general feature of any discrete variable system (e.g., Potts models, clock models, or lattice gauge theories).
The Finite-Size Scaling framework employed is universally applicable to second-order phase transitions. For systems in different universality classes, we predict that:
  1. 1.
    Monofractal universality classes (e.g., 3D Ising, $O(N)$ models with $N>1$, percolation): The spectral width $\Delta\alpha$ should collapse to zero under FSS with $q>0$, recovering the unique scaling dimension of each class. For example, the 3D Ising model should yield $\alpha\approx 0.964$ (from $\eta\approx 0.036$[13], while the 2D XY model should give $\alpha\approx 0.875$ (from $\eta\approx 0.25$).
  2. 2.
    Systems with quenched disorder (random field Ising, spin glasses, diluted models): Genuine broad multifractal spectra should persist, reflecting the spatial heterogeneity of local critical temperatures. Our RBIM results (Fig. 5) demonstrate the method’s sensitivity to this regime.
  3. 3.
    First-order transitions: The absence of scale invariance implies no genuine multifractal structure. Any observed broad spectra would be pure artifacts of finite size and should not exhibit systematic FSS collapse.
We caution that systems with continuously varying critical exponents (e.g., Kosterlitz-Thouless transitions, certain quasiperiodic systems) may require modified analysis. Additionally, the polynomial detrending order $\nu$ should be adapted to the analyticity of irrelevant operators in each universality class—higher-order corrections may necessitate $\nu\geq 2$.
The protocol is immediately applicable to experimental systems where MFDFA is routinely employed but the underlying universality class may be uncertain: magnetic materials near criticality, liquid crystals, ferroelectrics, and soft matter systems exhibiting critical-like scaling. The discriminatory power of $\Delta\alpha$ as an order parameter distinguishing “clean” versus “dirty” criticality provides a practical diagnostic for experimentalists.
Future extensions to continuous field theories (e.g., $\phi^{4}$ models in lattice field theory), higher-dimensional systems ($d>3$), and quantum critical points would establish the full scope of this methodology.
VII Conclusion
This work resolves a persistent contradiction in the literature: the apparent multifractal character of the 2D Ising model’s critical point, which violates Conformal Field Theory’s prediction of monofractal scaling. We demonstrate that this spurious multifractality is not an intrinsic property of the critical fixed point, but rather a finite-size artifact arising from the discreteness of spin variables in the negative moment regime ($q<0$). By restricting the analysis to positive moments ($q>0$) and implementing systematic Finite-Size Scaling analysis, we show that the singularity spectrum width collapses to zero ($\Delta\alpha\to 0$) in the thermodynamic limit, recovering the precise monofractal scaling dimension $\alpha\approx 0.875$ predicted by CFT. This provides strong support that MFDFA, when properly implemented, is consistent with key aspects of Renormalization Group principles, although further studies would be valuable to fully substantiate this connection.
The physical reinterpretation of MFDFA’s polynomial detrending as a phenomenological RG filter illuminates why this method succeeds. By suppressing analytic background fields associated with irrelevant operators, detrending isolates the singular critical fluctuations even in finite systems where correction-to-scaling terms are significant. This theoretical perspective suggests systematic protocols for improving exponent extraction in other universality classes by adapting the detrending order.
Our comparative analysis with the Random Bond Ising Model provides crucial validation. The RBIM exhibits a broad, robust multifractal spectrum ($\Delta\alpha\approx 0.23$) that persists under FSS, demonstrating that our protocol is sufficiently sensitive to detect genuine multifractality when physically present. This establishes MFDFA not merely as a numerical tool, but as a sensitive diagnostic capable of distinguishing “clean” monofractal criticality from “dirty” disorder-dominated complexity. The spectral width $\Delta\alpha$ emerges as an effective order parameter for this distinction.
The rigorous protocol established here enables MFDFA to function as a quantitative probe of universality classes in discrete lattice systems. By controlling finite-size effects through systematic FSS and choosing moment ranges to reflect the physical meaning of fluctuations, practitioners can reliably extract critical exponents and identify universality classes—even in experimental systems with unknown underlying Hamiltonians. The open-source implementation we provide ensures that this methodology is accessible to the broader computational physics community, fostering reproducible and reliable characterization of critical phenomena across disciplines and application domains.
Extensions of this framework offer promising directions for future research. Application to three-dimensional systems (3D Ising, $O(N)$ models) will quantitatively test our predictions for exponents in different universality classes. Adaptation to dynamical critical phenomena and quantum phase transitions will explore whether the RG-filter interpretation of detrending extends beyond equilibrium statistical mechanics. Systems with emergent scale invariance—including self-organized criticality and complex networks—represent another frontier where careful moment-range selection may illuminate genuine multifractal signatures. Most ambitiously, this work suggests a deeper connection between signal processing and RG theory: polynomial detrending, typically viewed as a technical necessity, may implement a systematic suppression of irrelevant operators, opening new avenues for extracting universal physics from finite-system measurements.